OpenStudy (anonymous):

Can I use logarithms to solve this? If so, how? 24000 + 3000y = 30000 * 1.04^(y-1)

OpenStudy (anonymous):

no, not with the 'y' there on the ground floor

OpenStudy (anonymous):

you can certianly cancel a bunch of stuff and get \[24+3y=30(1.04)^{y-1}\] or maybe even \[8+y=10(1.04)^{y-1}\] but then you are stuck

OpenStudy (anonymous):

okay, then how would i go about solving this word problem:

OpenStudy (anonymous):

this girl has two different job offers, job A starting salary is 27000 with an increase of 3000 each year after the first, job B has a starting salary of 30000 with a 4% increase each year after the first year

OpenStudy (anonymous):

in how many years will job A's salary be higher than job B's? The answer is obviously two years, but how would you do it mathematically?

OpenStudy (anonymous):

when will they be equal?

OpenStudy (anonymous):

trial and error

OpenStudy (anonymous):

somewhere betwen 2 and 3 years

OpenStudy (anonymous):

I made two equations then set them equal to each other, and then whatever answer i got i would round up. A = 27000 + 3000(y-1) B = 30000 * 1.04^(y-1)

OpenStudy (anonymous):

It's on an open response so i cant just put "guess and check"

OpenStudy (anonymous):

A 27,000 30,000 33,000 36,000 B 30,000 31,200 32448, 33745.92

OpenStudy (anonymous):

I understand that, but I can't just put a table of the values, I need to show it algebraically. Is there any way to do that?